# On the cardinality of measures of maximal relative entropy for smooth skew products

### v2

@@ 4.1. Petersen-Quas-Shin construction
$h_{\mu^{3}}(\sigma_{X})$ $=\lim_{n\to\infty}h_{\mu^{3}_{n}}(\sigma_{X_{n}})$
$\geq\lim_{n\to\infty}\left[\frac{h_{\mu^{1}_{n}}(\sigma_{X_{n}})+h_{\mu^{2}_{n}}(\sigma_{X_{n}})}{2}+\int_{S_{n}}\Xi_{n}(z)\,\eta_{n}(\mathrm{d}z)\right]$
$\geq\lim_{n\to\infty}\left[\frac{h_{\mu^{1}_{n}}(\sigma_{X_{n}})+h_{\mu^{2}_{n}}(\sigma_{X_{n}})}{2}+\int_{S_{n}}\psi\left(\frac{A_{n}^{j_{0}}+B_{n}^{j_{0}}}{2}\right)-\frac{1}{2}\psi(A_{n}^{j_{0}})-\frac{1}{2}\psi(B_{n}^{j_{0}})\mathrm{d}\eta_{n}\right]$
$>\frac{h_{\mu^{1}}(\sigma_{X})+h_{\mu^{2}}(\sigma_{X})}{2}=h_{\mu^{1}}(\sigma_{X}).$
@@
Before proving Theorem A, we quote two results from the literature and prove a lemma. The first is a consequence of [54, Theorems 4.4 and 5.1] and the second is a result proved in [45].
Lemma 5.1 ([54, Lemma 3.1, Theorems 4.4 and 5.1]).
Let $(\Sigma^{\prime}_{b},\sigma_{b})$ be a subshift of finite type, $\Lambda$ be topological space, $\theta:\Lambda\to\Lambda$ a homeomorphism, $\pi:\Sigma\to\Lambda$ a surjective finite-to-one Hölder continuous map. Then,
  • (i)
    for each ergodic $\mathbb{P}\in\mathcal{M}_{e}(\theta)$ there exists $\widehat{\mathbb{P}}\in\mathcal{M}_{e}(\sigma)$ such that $(\pi)_{*}\widehat{\mathbb{P}}$$\mathbb{P}$;
  • (ii)
    there exists only finitely many ergodic probability measures $\widehat{\mathbb{P}}_{1},\ldots,\widehat{\mathbb{P}}_{k}\in\mathcal{M}_{e}(\theta)$, such that $(\pi)_{*}\widehat{\mathbb{P}}_{i}=\mathbb{P}$; and
  • (iii)
    there exists $N_{\mathbb{P}}\in\mathbb{N}$ such that $\pi^{-1}(\omega)=N_{\mathbb{P}}$ for $\mathbb{P}$-almost every $\omega\in\Lambda$.
We note that [45, Proposition 13.2] treats a slightly different setting, but the proof applies in our context because it uses only the finite-to-one coding properties (see also [10, Thm. 0.2] and [19, Thm. 10.5]).
Lemma 5.2 ([45, Proposition 13.2 and Step 2 of its proof]).
@@
Let $\mathbb{P}$ be a $\theta$-invariant ergodic measure. By Lemma 5.3, it suffices to show that $\mathcal{M}_{\max}^{\mathcal{X},\chi}(\Theta\mid\mathbb{P})$ is countable for each homoclinic class $\mathcal{X}$ and each $\chi>0$.
Fix $\chi>0$ and a Borel homoclinic class $\mathcal{X}$. Let $\Sigma_{s}$ be an irreducible countable-state topological Markov shift, $\Sigma_{b}$ a subshift of finite type, and let $\pi_{s}:\Sigma_{s}\to\mathcal{X}$ and $\pi_{b}:\Sigma_{b}\to\Omega$ be the Hölder maps given by Theorem 3.5. Recall that:
@@
    $\pi_{s}[\Sigma_{s}^{\#}]$ has full measure for every $\chi$-hyperbolic $\mu\in\mathcal{M}_{e}(\Theta)$ with $\mu(\mathcal{X})=1$; moreover $\pi_{s}:\Sigma_{s}^{\#}\to\mathcal{X}$ is finite-to-one, and $\pi_{b}:\Sigma_{b}\to NW(\theta)$ is a surjective finite-to-one map;
@@
    and $\Sigma_{b}^{\prime}:=\mathrm{proj}_{\Sigma_{b}}(\Sigma_{s})$ is an irreducible subshift of finite type.
@@
Thus $\widehat{\mathbb{P}}:=(\mathrm{proj}_{\Sigma_{b}})_{*}\widehat{\mu}$ is ergodic as a factor of $\widehat{\mu}$, and is a $\sigma_{b}$-invariant $\pi_{b}$-lift of $\mathbb{P}$ supported on $\Sigma_{b}^{\prime}$.
Claim. For the factor map $\mathrm{proj}_{\Sigma_{s}}:(\Sigma_{s},\sigma_{s})\to(\Sigma_{b}^{\prime},\sigma_{b})$, $\widehat{\mu}$ is a measure of maximal $^\widehat{\mathbb{P}}$-relative entropy. When the context is clear, we simply say that $\widehat{\mu}$ is a measure of maximal entropy relative to $\widehat{P}$.
@@
By Lemmas 5.1(iii) and 5.2(iii), $\pi_{b}$ is $N_{\mathbb{P}}$-to-one on a set of $\mathbb{P}$-full measure and $\pi_{s}$ is $N_{\eta}$-to-one on a set of $\eta$-full measure; since finite extensions preserve entropy, we have
@@
For each $\mu\in\mathcal{M}_{\max}^{\mathcal{X},\chi}(\Theta\mid\mathbb{P})$ we have thus associated a measure $\widehat{\mu}\in\mathcal{M}_{e}(\sigma_{s})$ such that $\widehat{\mu}$ is a measure of maximal entropy relative to $\widehat{\mathbb{P}}:=(\mathrm{proj}_{\Sigma_{b}})_{*}\widehat{\mu}$ and
$(\pi_{b}\circ\mathrm{proj}_{\Sigma_{b}})_{*}\widehat{\mu}$$\mathbb{P}$.
Define
$\mathcal{M}_{\max}(\Sigma_{s}\mid\mathbb{P}):=\left\{\widehat{\mu}\in\mathcal{M}_{e}(\sigma_{s})\ \middle|\ \begin{aligned} &\widehat{\mu}\ \text{has maximal entropy relative to }(\mathrm{proj}_{\Sigma_{b}})_{*}\widehat{\mu}\\ &\text{and }(\pi_{b}\circ\mathrm{proj}_{\Sigma_{b}})_{*}\widehat{\mu}$$\mathbb{P}\end{aligned}\right\}$.
@@
By Lemma 5.1, there exist only finitely many ergodic $\sigma_{b}$-invariant probability measures $\widehat{\mathbb{P}}_{1},\dots,\widehat{\mathbb{P}}_{k}$ supported on $\Sigma_{b}^{\prime}$ such that $(\pi_{b})_{*}\widehat{\mathbb{P}}_{i}=\mathbb{P}$. Hence, using the injectivity of (17),
$\#\,\mathcal{M}_{\max}^{\mathcal{X},\chi}$ $(\Theta\mid\mathbb{P})\leq\#\,\mathcal{M}_{\max}(\Sigma_{s}\mid\mathbb{P})$
$=\#\left(\bigcup_{i=1}^{k}\left\{\widehat{\mu}\in\mathcal{M}_{e}(\sigma_{s});\ \widehat{\mu}\ \text{has maximal entropy relative to }\widehat{\mathbb{P}}_{i}\right\}\right).$ (18)
From Corollary 4.5, with $X=\Sigma_{s}$, $Y=\Sigma_{b}^{\prime}=\mathrm{proj}_{\Sigma_{b}}(\Sigma_{s})$, $\pi=\mathrm{proj}_{\Sigma_{b}}$ and $\nu=\widehat{\mathbb{P}}_{i}$ it follows that, for each $i\in\{1,\dots,k\}$, the set
$\left\{\widehat{\mu}\in\mathcal{M}_{e}(\sigma_{s}):\,\widehat{\mu}\ \text{has maximal entropy relative to }\widehat{\mathbb{P}}_{i}\right\}$
is at most countable. Therefore from (18) we obtain that $\mathcal{M}_{\max}^{\mathcal{X},\chi}(\Theta\mid\mathbb{P})$ is countable, which completes the proof. ∎
To prove Corollary B, we recall the following theorem for skew products, which refines Ruelle’s inequality. We state a convenient weak version of the result presented in [25] that suffices for our purposes.
Theorem 5.4 ([25, Chapter 5, Theorem 3.1.1]).
@@
with $\lambda_{+}<0<\lambda_{+}$. From Theorem 5.4 we have that
@@
By construction $\gamma_{\ell}\in\mathcal{G}_{r_{\ell}}$. Therefore, there exists $g_{\ell}:J_{r_{\ell}}\to\mathbb{R}$ such that $\gamma_{\ell}(t)=(t,g_{\ell}(t))$. Thus, $\gamma_{\ell}(J_{r_{\ell}})=\{(t,g_{\ell}(t));t\in J_{r_{\ell}}\}\subset S_{r_{\ell}}$, and from (22) we obtain that $(\mathcal{T}_{\ell})_{\ell\in\mathbb{N}}$ is a nested sequence of non-empty compact sets, hence
$\bigcap_{\ell=0}^{\infty}\mathcal{T}_{\ell}\neq\emptyset.$
@@
By Step 2, $X_{\omega}\neq\emptyset$, and the coding map
@@ Acknowledgements
MMC acknowledges useful early discussions with Anthony Quas and Yuri Lima. GF and MMC are supported by an Australian Research Council Laureate Fellowship (FL230100088).

### v3

@@ 4.1. Petersen-Quas-Shin construction
$h_{\mu^{3}}(\sigma_{X}\mid\nu)$ $=\lim_{n\to\infty}h_{\mu^{3}_{n}}(\sigma_{X_{n}}\mid\nu)$
$\geq\lim_{n\to\infty}\left[\frac{h_{\mu^{1}_{n}}(\sigma_{X_{n}}\mid\nu)+h_{\mu^{2}_{n}}(\sigma_{X_{n}})}{2}+\int_{S_{n}^{(a)}}\Xi_{n}(z)\,\eta_{n}(\mathrm{d}z)\right]$
$\geq\lim_{n\to\infty}\left[\frac{h_{\mu^{1}_{n}}(\sigma_{X_{n}})+h_{\mu^{2}_{n}}(\sigma_{X_{n}})}{2}+\int_{S_{n}^{(a)}}\psi\left(\frac{A_{n}^{b}+B_{n}^{b}}{2}\right)-\frac{1}{2}\psi(A_{n}^{b})-\frac{1}{2}\psi(B_{n}^{b})\mathrm{d}\eta_{n}\right]$
$>\frac{h_{\mu^{1}}(\sigma_{X}\mid\nu)+h_{\mu^{2}}(\sigma_{X}\mid\nu)}{2}=h_{\mu^{1}}(\sigma_{X}\mid\nu).$
@@
Before proving Theorem A, we recall a result from the literature and prove a lemma. The proof of the Lemma below was sketched in Step 5 of the proof of Theorem 3.6.
Lemma 5.1 ([Sarig2013, Proposition 13.2 and Step 2 of its proof]).
@@
Let $\mathbb{P}$ be a $\theta$-invariant ergodic probability measure with full support. By Lemma 5.2, it suffices to show that $\mathcal{M}_{\max}^{\mathcal{X},\chi}(\Theta\mid\mathbb{P})$ is countable for each homoclinic class $\mathcal{X}$ and each $\chi>0$.
Fix $\chi>0$ and a Borel homoclinic class $\mathcal{X}$. Let $\Sigma_{s}$ be an irreducible locally compact topological Markov shift, $\Sigma_{b}$ a subshift of finite type, and let $\pi_{s}:\Sigma_{s}\to\mathcal{X}$ and $\pi_{b}:\Sigma_{b}\to\Omega$ be the Hölder maps given by Theorem 3.6. Recall that:
@@
    $\pi_{s}[\Sigma_{s}^{\#}]$ has full measure for every $\chi$-hyperbolic $\mu\in\mathcal{M}_{e}(\Theta)$ with $\mu(\mathcal{X})=1$; moreover $\pi_{s}:\Sigma_{s}^{\#}\to\mathcal{X}$ is finite-to-one, and $\pi_{b}:\Sigma_{b}\to\Omega$ is a surjective finite-to-one map;
@@
    with $\widehat{\mathbb{P}}[\mathrm{proj}_{\Sigma_{b}}(\Sigma_{s})]=1$, where $\widehat{\mathbb{P}}$ unique measure in $M_{e}(\sigma_{b})$ such that $(\pi_{b})_{*}\widehat{\mathbb{P}}$$\mathbb{P}$.
@@
Thus $(\mathrm{proj}_{\Sigma_{b}})_{*}\widehat{\mu}$$\widehat{\mathbb{P}}$, since the lift of $\mathbb{P}$ via $\pi_{b}$ is unique (recall that $\mathbb{P}$ is assumed to be fully supported on $\Omega$).
Claim. For the factor map $\mathrm{proj}_{\Sigma_{b}}:(\Sigma_{s},\sigma_{s})\to(\Sigma_{b},\sigma_{b})$, $\widehat{\mu}$ is a measure of maximal $^\widehat{\mathbb{P}}$-relative entropy.
@@
By Lemma 5.1(iii) $\pi_{s}$ is $N_{\eta}$-to-one on a set of $\eta$-full measure. Moreover, recall that since $\mathbb{P}$ is fully supported on $\Omega$, $\pi_{b}$ is $\mathbb{P}$-a.s. one-to-one [BowenBook, 3.18]. Using that finite extensions preserve entropy, we have
@@
For each $\mu\in\mathcal{M}_{\max}^{\mathcal{X},\chi}(\Theta\mid\mathbb{P})$ we have thus associated a measure $\widehat{\mu}\in\mathcal{M}_{e}(\sigma_{s})$ such that $\widehat{\mu}$ is a measure of maximal $^\widehat{\mathbb{P}}$-relative entropy. Moreover, $(\pi_{b}\circ\mathrm{proj}_{\Sigma_{b}})_{*}\widehat{\mu}$$\mathbb{P}$. Define
$\mathcal{M}_{\max}(\Sigma_{s}\mid\mathbb{P})$ $:=\left\{\widehat{\mu}\in\mathcal{M}_{e}(\sigma_{s})\ ;\ \widehat{\mu}\ \text{has maximal }\widehat{\mathbb{P}}\text{-relative entropy}\right\}$
@@
Hence, using the injectivity of (32), we obtain that
$\#\,\mathcal{M}_{\max}^{\mathcal{X},\chi}$ $(\Theta\mid\mathbb{P})\leq\#\left(\left\{\widehat{\mu}\in\mathcal{M}_{e}(\sigma_{s});\ \widehat{\mu}\ \text{has maximal $\widehat{\mathbb{P}}$-relative entropy}\right\}\right).$ (33)
Observe that since $\mathbb{P}$ is fully supported on $\Omega$, then $\widehat{\mathbb{P}}$ is fully supported on $\Sigma_{b}$ (as a consequence of [BowenBook, Theorem 3.19 and Proposition 3.19]). From Corollary 4.5, with $X=\Sigma_{s}$, $Y=\Sigma_{b}=\mathrm{proj}_{\Sigma_{b}}(\Sigma_{s})$, $\pi=\mathrm{proj}_{\Sigma_{b}}$ and $\nu=\widehat{\mathbb{P}}$ it follows that the set
$\left\{\widehat{\mu}\in\mathcal{M}_{e}(\sigma_{s}):\,\widehat{\mu}\ \text{has maximal $\widehat{\mathbb{P}}$-relative entropy}\right\}$
is at most countable. Therefore from (33) we obtain that the set $\mathcal{M}_{\max}^{\mathcal{X},\chi}(\Theta\mid\mathbb{P})$ is countable, which completes the proof. ∎
To prove Corollary B, we recall the following theorem for skew products, which refines Ruelle’s inequality. We state a convenient weak version of the result presented in [KatokKifer] that suffices for our purposes.
Theorem 5.3 ([KatokKifer, Chapter 5, Theorem 3.1.1]).
@@
with $\lambda_{-}<0<\lambda_{+}$. From Theorem 5.3 we have that
@@
By construction $\gamma_{\ell}\in\mathcal{G}_{r_{\ell}}$. Therefore, there exists $g_{\ell}:J_{r_{\ell}}\to\mathbb{R}$ such that $\gamma_{\ell}(t)=(t,g_{\ell}(t))$. Thus, $\gamma_{\ell}(J_{r_{\ell}})=\{(t,g_{\ell}(t));t\in J_{r_{\ell}}\}\subset S_{r_{\ell}}$, and from (37) we obtain that $(\mathcal{T}_{\ell})_{\ell\in\mathbb{N}}$ is a nested sequence of nonempty compact sets, hence
$\bigcap_{\ell=0}^{\infty}\mathcal{T}_{\ell}\neq\varnothing.$
@@
By Step 2, $X_{\omega}\neq\varnothing$, and the coding map
@@ Acknowledgements
The authors acknowledge Yuri Lima for pointing out an inconsistency in the proof of Theorem 3.6 in an earlier version, and Anthony Quas for valuable discussions and for suggesting an improved approach for the proof of Theorem 4.4. GF and MMC are supported by an Australian Research Council Laureate Fellowship (FL230100088).